CAP Workshop 1991

Proceedings

of the

Second Fujitsu-ANU CAP Workshop

Australian National University

November 28-29, 1991

Appeared as Proceedings of the Second Fujitsu-ANU CAP Workshop (edited by R. P. Brent), Computer Science Department, ANU, November 1991, 254 pp. Only the title page, preface, acknowledgements, program and index are given here.

All Rights Reserved. Copyright ◯^{c} 1991 Australian National University and Fujitsu Laboratories Limited.

rpb129pr typeset using TEX

Preface

The CAP Workshop 1991 is the second in a series of workshops supported by Fujitsu under the auspices of the Fujitsu-ANU joint research agreement. The first workshop was held at Fujitsu Research Laboratories, Kawasaki, Japan, in November 1990. The second workshop is being held at the Australian National University, Canberra, Australia, in November 1991.

The organisers would like to thank all those who have contributed to the success of this workshop –

• Fujitsu Ltd., Fujitsu Research Laboratories, and Fujitsu Australia for their support of the Fujitsu-ANU joint research agreement;

• the Centre for Information Science Research at ANU for its support and coordinating role;

• members of the Computer-Based Systems Laboratory at Fujitsu Research Laboratories and the ANU CAP Team for their technical assistance;

• members of the Advisory Board, Program Committee and Local Arrangements Committee;

• Michelle Moravec for assistance with production of the Proceedings;

• contributors to the panel session for their insights;

• and not least the authors of the papers for the time and effort involved in preparing their papers.

CAP Workshop Organisation

Advisory Board and Program Committee

Richard Brent Australian National University Yoshihiro Kosaka Fujitsu Ltd. Robin Stanton Australian National University Junichi Tanahashi Fujitsu Laboratories Ltd.

Local Arrangements and Coordination

Sally Begg Australian National University David Hawking Australian National University Mitsuo Ishii Fujitsu Laboratories Ltd. Takeshi Okuzumi Fujitsu Ltd. Peter Price Australian National University Hiroyuki Sato Fujitsu Laboratories Ltd. Robin Stanton Australian National University Peiyi Tang Australian National University

Proceedings Editor

Richard Brent Australian National University

Program – Day 1

Thursday, 28 November 1991

8:30 - 9:00 Registration
9:00 - 9:30 Opening Session
9:30 - 10:30 Paper Session 1: Hardware and Systems
A.  Low-Latency Message Communication Support for the AP 1000

T. Shimizu, T. Horie and H. Ishihata

B.  All-to-All Personalized Communication on Wrap-around Mesh

T. Horie and K. Hayashi

10:30 - 11:00 Break
11:00 - 12:30 Paper Session 2: Software Tools and Compilers
C.  Developing Monitoring and Debugging Tools

for the AP 1000 Array Multiprocessor

C. W. Johnson, P. B. Thistlewaite, D. Walsh and M. Zellner

D.  Parallel Loop Code Generation for the AP 1000

Gavin Michael and Peiyi Tang

E.  Data Parallelism on MIMD Machines

T. Bossomaier

12:30 - 13:30 Lunch
13:30 - 15:00 Paper Session 3: Linear Algebra
F.  A Parallel Direct Solution Method of Sparse Linear Systems

and its Evaluation on AP 1000 K. Motegi and S. Watanabe

G.  The LINPACK Benchmark on the AP 1000: Preliminary Report

R. P. Brent

H.  The Implementation of BLAS level 3 on the AP 1000: Preliminary Report

P. E. Strazdins and R. P. Brent

J.  Singular Value Computation on the AP 1000

A. Czezowski and P. E. Strazdins

15:00 - 15:30 Break
15:30 - 17:00 Paper Session 4: Applications
K.  Application of Parallel Computer to Protein Engineering

H. Iwama, S. Kawakita, M. Saito, H. Sato, Y. Tanaka, S. Tsutsumi and H. Yoshijima

L.  Parallelization Strategies for MD Simulations on the AP 1000

D. Brown and J. H. R. Clarke

M.  Parallelization of the Monte Carlo Codes MCACE for Shielding Analysis

and Measurement of Parallel Efficiency on AP 1000 with 64 Cell Processors M. Takano, F. Masukawa, Y. Naito, A. Kawazoe and M. Okuda

Program – Day 2

Friday, 29 November 1991

9:00 - 10:30 Paper Session 5: Hardware, Systems and Visualization
N.  Performance Evaluation of the AP 1000

H. Ishihata, T. Horie, T. Shimizu and S. Kato

P.  Connecting AP 1000 with mainframe:

Application to KEK Experimental HEP computations S. Ichikawa, N. Ishida, S. Yamaki, T. Matsuura, A. Manabe, K. Amako, H. Fujii, S. Ichii, R. Itoh, A. Miyamoto, Y. Takaiwa and Y. Watase

Q.  Scientific Visualization Algorithms on the AP 1000

P. Mackerras

10:30 - 11:00 Break
11:00 - 13:00 Paper Session 6: File Systems, Software and Applications
R.  Implementation and Performance of the Acacia File System

B. M. Broom

S.  High Speed Search of Large Text Bases on the Fujitsu AP 1000

D. Hawking

T.  Implementation of C-Linda for the AP 1000

R. Cohen and B. Molinari

U.  PARACS: A Highly Parallel Circuit Simulator

J. Niitsuma, K. Teramae, T. Kage, S. Shimogori and Y. Izuta

13:00 - 14:00 Lunch
14:00 - 15:30 Paper Session 7: Applications
V.  Particle Simulation using Monte Carlo Method

M. Okuda, M. Fujisaki, K. Watanable, A. Kawazoe, M. Yokokawa, H. Yamamoto, H. Kaburaki, S. Inawashiro and F. Matubara

W.  QCD Lattice Monte Carlo Simulations on Cell Array Processor AP 1000

K. Akemi, Ph. de Forcrand, M. Fujisaki, T. Hashimoto,

H. C. Hege, S. Hioki, J. Makino, O. Miyamura, A. Nakamura, M. Okuda, I. O. Stamatescu, Y. Tago and T. Takaishi

X.  Combinatorial Optimisation on the AP 1000

K. Malysiak and B. D. McKay

15:30 - 16:00 Break
16:00 - 17:00 Panel Discussion

The Future of Distributed-Memory Parallel Computers

R. Brent (chair), T. Gojobori, M. Ishii,

H. Kasahara, Y. Muraoka, R. Stanton and S. Tomita

17:00 - 17:30 Closing Session

Bon Voyage !

Author Index

Author Paper(s)  Author Paper(s)
Akemi  W  Matsuura P
Amako  P  Matubara V
Bossomaier E  McKay  X
Brent  G, H  Michael  D
Broom  R  Miyamoto P
Brown  L  Miyamura W
Clarke  L  Molinari  T
Cohen  T  Motegi  F
Czezowski J  Naito  M
Forcrand  W  Nakamura W
Fujii  P  Niitsuma  U
Fujisaki  V, W  Okuda  M, V, W
Hashimoto W  Saito  K
Hawking  S  Sato  K
Hayashi  B  Shimizu  A, N
Hege  W  Shimogori U
Hioki  W  Stamatescu W
Horie  A, B, N  Strazdins  H, J
Ichii  P  Tago  W
Ichikawa  P  Takaishi  W
Inawashiro V  Takaiwa  P
Ishida  P  Takano  M
Ishihata  A, N  Tanaka  K
Itoh  P  Tang  D
Iwama  K  Teramae  U
Izuta  U  Thistlewaite C
Johnson  C  Tsutsumi  K
Kaburaki  V  Walsh  C
Kage  U  Watanabe F
Kato  N  Watanable V
Kawakita  K  Watase  P
Kawazoe  M, V  Yamaki  P
Mackerras Q  Yamamoto V
Makino  W  Yokokawa V
Malysiak  X  Yoshijima K
Manabe  P  Zellner  C
Masukawa M